# Anubis: Towards Reliable Cloud AI Infrastructure via Proactive Validation

### v1

@@ 5.2 Benchmark Selection Evaluation
    If the Selector misses defects and there exists incident during job running, a new benchmark selection will be performed based on updated node statuses. The job will be pushed to the queue rear and continue where it lefts off.
@@
The average node utilization results are presented in Figure 8. Anubis Selector with benchmark selection algorithm achieves a high cluster utilization of $90.70\%$, improving the no validation baseline by 4.81×4.81$\times$4.81 × and the full set baseline by 1.09×1.09$\times$1.09 ×. As depicted in Table 4, the Selector reduces $92.07\%$ validation time cost per node compared to the full set baseline, which is irrelevant to simulation duration since validation is performed regularly in production. In terms of MTBI, the Selector achieves a high $262.05$ hours, enhancing the no validation baseline by 22.61×22.61$\times$22.61 × and the full set baseline by $1\times1$. The Selector experiences an average of 4.79 incidents per node, 0.16 higher than full set baseline. Despite more incidents, it benefits from reduced down time due to selective validation, resulting in longer up time and better MTBI.
@@
Policy Absence Full Set Anubis Selector
@@ 5.4 Deployment in Cloud
We have deployed Anubis in Azure production environment for two years, where it acts as a quality gate to identify most defective nodes before delivering GPU clusters to public cloud services. We collect the benchmark dataset during the cluster build-out phase from 24k$+$ A100 GPUs (3k$+$ VMs) in 90 days and demonstrate its practicality.
Benchmark Effectiveness
Repeatability at Scale
@@ 6 Related Works
Performance benchmarking is used to measure which computer system can complete tasks more quickly across different types. For traditional CPU-based systems, SPEC [20] and LINPACK [22] are widely-used to rank supercomputers. For AI systems, particularly GPU-based, DeepBench [56] compares typical operations in deep learning, while DAWNBench [18] and MLPerf [42] consider end-to-end training or inference performance and measure time to convergence as metrics.
@@ 7 Conclusion
Inherent redundancies in cloud AI infrastructure can give rise to a unique degradation pattern for AI workloads, leading to diminished reliability. In this paper, we delved into incidents associated with this concealed degradation pattern and addressed it through the design of a proactive validation system, Anubis. The system comprises a comprehensive benchmark suite, a Selector designed to efficiently trade-off validation time and coverage, and a Validator which executes benchmarks and filters defects with clear-cut criteria. The system also iterates and evolves with the rich data gathered by itself. Our evaluation in testbed and simulation demonstrates its ability to increase the MTBI by 22.61×22.61$\times$22.61 ×. Moreover, Anubis has showcased its practicality in a real-world cloud deployment.
Appendices
Appendix A Optimizations in Networking Validation
For the distributed micro-benchmark which performs pairwise scans for RDMA verbs to ensure the networking quality of InfiniBand or RoCEv2 during large-scale AI training, we propose the following two algorithms to reduce validation time:
Full Scan in $\mathcal{O}(n)$
The pairwise scans for all nodes is to ensure that the networking bandwidth between any two nodes are expected so that AI workloads with all-to-all traffic (e.g., Mixture-of-Experts models) won’t have regression. The key idea is to simultaneously run network tests for all nodes, which can avoid traffic collision by leveraging the Clos network [16] property. The problem can be formulated as follows. Given a network with ${N}$ NICs, suppose ${N}\in 2\leavevmode\nobreak\ \mathbb{N}$ and pairs are symmetric, schedule all possible $\frac{{N}\cdot({N}-1)}{2}$ NIC pairs into ${N}-1$ rounds. In each round, exactly $\frac{{N}}{2}$ pairs will run pairwise benchmarks (e.g., GPU-direct RDMA write [55], NCCL/RCCL [46, 6] 2-node all-reduce) without NIC intersection. We address this problem by leveraging the circle method in round-robin tournaments [34] as the scheduling algorithm.
Quick Scan in $\mathcal{O}(1)$
To eliminate the increasing time required for large-scale networking validation, we further propose topology-aware networking validation, which scans network links within a fixed duration, irrespective of the network scale or node number. We select node pairs such that the distance (number of hops) between each pair is 2-hop (nodes under the same ToR switch), 4-hop (nodes under the same aggregation switch), 6-hop (nodes across the core switch), and so on. For each hop number, we include all nodes without node overlap, ensuring that every node is included once and only once for the given hop. For a $k$-tier fat-tree network topology, only $k$ rounds are needed for any scale. In each round, we run the selected networking benchmarks for each hop number.
Appendix B Benchmark Parameter Searching
For end-to-end benchmarks, since model structures and hyper-parameters like batch size and sequence length have already been determined based on representative workloads, the search focuses on the number of warmup steps $w$ and measurement steps $n$. It aims to minimize the total step number, i.e., benchmark running time, while maximizing the repeatability across nodes. This process is performed offline.

### v2

@@ 5.2 Benchmark Selection Evaluation
    If the Selector fails to predict defects that cause incidents during job running, a new benchmark selection will be performed based on updated node statuses. The job will be pushed to the queue rear and continue where it lefts off.
@@
Policy Absence Full Set SuperBench Selector
@@
The average node utilization results are presented in Figure 8. SuperBench Selector with benchmark selection algorithm achieves a high cluster utilization of $90.70\%$, improving the no validation baseline by 4.81×4.81$\times$4.81 × and the full set baseline by 1.09×1.09$\times$1.09 ×. As depicted in Table 4, the Selector reduces $92.07\%$ validation time cost per node compared to the full set baseline, which is irrelevant to simulation duration since validation is performed regularly in production. In terms of MTBI, the Selector achieves a high $262.05$ hours, enhancing the no validation baseline by 22.61×22.61$\times$22.61 × and the full set baseline by $1\times1$. The Selector experiences an average of 4.79 incidents per node, 0.16 higher than full set baseline. Despite more incidents, it benefits from reduced down time due to selective validation, resulting in longer up time and better MTBI.
@@ 5.4 Deployment in Cloud
We have deployed SuperBench in Azure production environment for two years, where it acts as a quality gate to identify most defective nodes before delivering GPU clusters to public cloud services. We collect the benchmark dataset during the cluster build-out phase from 24k$+$ A100 GPUs (3k$+$ VMs) in 90 days and demonstrate its practicality.
Benchmark Effectiveness
Repeatability at Scale
@@ 6 Related Works
Performance benchmarking is used to measure which computer system can complete tasks more quickly across different types. For traditional CPU-based systems, SPEC [21] and LINPACK [23] are widely-used to rank supercomputers. For AI systems, particularly GPU-based, benchmarking also shares the same goal of measuring which AI system is faster. DeepBench [57] compares typical operations in deep learning, while DAWNBench [18] and MLPerf [43] consider end-to-end training or inference performance and measure time to convergence as metrics.
@@ 7 Conclusion
Inherent redundancies in cloud AI infrastructure can give rise to a unique degradation pattern for AI workloads, leading to diminished reliability. In this paper, we delved into incidents associated with this concealed degradation pattern and addressed it through the design of a proactive validation system, SuperBench. The system comprises a comprehensive benchmark suite, a Selector designed to efficiently trade-off validation time and coverage, and a Validator which executes benchmarks and filters defects with clear-cut criteria. The system also iterates and evolves with the rich data gathered by itself. Our evaluation in testbed and simulation demonstrates its ability to increase the MTBI by 22.61×22.61$\times$22.61 ×. Moreover, SuperBench has showcased its practicality in a real-world cloud deployment.
Acknowledgments
We sincerely thank our colleagues from Systems and Networking Research Group at Microsoft Research Asia and Azure HPC/AI team for their early and valuable feedback on this work. We would also like to thank the anonymous reviewers for their detailed and valuable feedback on this paper.